OpenStudy (anonymous):

Can anyone help me with this. find the following antiderivative of ∫(sin^4 x) (cos^3 x) dx.

OpenStudy (anonymous):

sin^2 x + cos ^2 x = 1

OpenStudy (anonymous):

lop of one of the sin factors, then use the half angle forumals.

OpenStudy (anonymous):

can I let u=cos x, if I want to use substitution ?

OpenStudy (anonymous):

after using the half angle forumals, you can let sin x equal u

OpenStudy (anonymous):

yea, use u = sin, integrand becomes (u^4 )(cos^2 x) = u^4( 1- u^2 ) = u^4-u^6 du

OpenStudy (anonymous):

ok. I got the idea. means I have to change cos^2 x = 1- sin^2 x right ? can I do this ? (sin^4 x) ( cos^3 x )=(sin^4 x ) (1- sin^2 x ) cos x dx then I let u= sin x. I'll get u^2 (1- u^2 ) du. is it right ?

OpenStudy (anonymous):

right with the sub u=sin x then you lose one of the cos x fomr the integrans leaving a cos^2 x which equals, because of the sub ( 1 - u^2)

OpenStudy (anonymous):

so the integral becomes u^4 ( 1 - u ^ 2 ) = u^4 - u^6 I put it in a symbolic integrator and go a different answer but numerically it works.

OpenStudy (anonymous):

ok. I got it. thanks a lot, roachie.

OpenStudy (anonymous):

yup